Suppose that in 1991 the total output in a single-good economy was 13,689 chickens. Also suppose that in 1991 each chicken was priced at $19. Finally, assume that in 1995 the price per chicken was $11 and that 19,561 were purchased.

For each of the following parts, assume that 1991 is the base year.

Note: Please round your answers to two decimal places.

a) Determine the GDP Deflator for 1995
1
Expert's answer
2020-03-09T09:39:52-0400

"GDP price deflator=(\frac{nominalGDP}{realGDP})∗100"

Nominal GDP is the current market prices while Real GDP is the market prices of the base year

"Nominal GDP =11 \times 19561" =215171

"Real GDP=19 \times 13689" =260091

GDP price deflator for "1995=215171/260091 \times 100" =82.73